On January 4,2021, Tamara Bakery paid ₱30 million for 1 million shares of jade Company ordinary shares. The investment represents a 20% interest in the net assets of jade and gave Tamara the ability to exercise significant influence over Jade’s operations. Tamara received dividends of ₱1.00 per share on December 15,2021 and jade reported net income of ₱8 million for the year ended December 31,2021. The market value of Jade’s ordinary shares at December 31,2021 was ₱32 per share, on the purchase date, the book value of Jade’s net assets was ₱120 million and the fair market value of Jade’s depreciable assets with an average remaining useful life of six years exceeded their book value by ₱6 million. The remainder of the excess of the cost of the investment over the book value of net assets purchased was attributable to goodwill.


What amount of investment revenue should Tamara report on its income statement for the year ended December 31, 2016, under the fair value method?


1
Expert's answer
2022-03-17T13:57:45-0400

"Investment revenue = dividends received + proportion of the capital gain in change of value\n\n= 1000000 + 0.2*2*1000000\n\n= P1000000 + P400000 = P1400000"